OpenStudy (anonymous):

Solve 8(4q-5)-7q=20q-10 in steps?

OpenStudy (anonymous):

Expand: \[32q-40-7q=20q-10\] Simplify: \[25q-40=20q-10\] \[5q=30\] \[q=6\]

OpenStudy (anonymous):

Thanks again :)